# Sampling-based Model Predictive Control Leveraging Parallelizable Physics Simulations

### v2

@@ Abstract
Code and videos: https://sites.google.com/view/mppi-isaac/
@@ I Introduction
As robots become increasingly integrated into our daily lives, their ability to navigate and interact with the environment is becoming more important than ever. From collision avoidance to moving obstacles out of the way to pick up some objects, robots must be able to plan their motions while accounting for contact with their surroundings. At the same time, robotic platforms are becoming more and more complex, and include many Degrees Of Freedom (DOF) in order to achieve agile and dexterous movements. All this poses many challenging problems to motion planners, such as collision-free navigation in complex and dynamic environments, high DOF mobile manipulation, contact-rich tasks such as picking and pushing, and in-hand manipulation.
Solutions to these challenges exist but are often specialized and not easily transferable to different scenarios. Learning-based approaches, for example, can leverage physics simulators to train policies for complex tasks but require extensive training and resources. For instance, [1] took years to develop, utilizing 6144 CPU cores and 50 hours of training to learn a policy for in-hand cube manipulation.
Figure 1: Scheme of the proposed method using IsaacGym as the dynamic model for MPPI. At each time step, IsaacGym is reset to the current world’s state $q$, and random input sequences ${V}$ are applied for the horizon $T$, to every environment. The resulting rolled-out trajectories are used by MPPI to approximate the optimal control $u_{0}^{*}$ given a cost function $C$.
On the other hand, model-based approaches like Model Predictive Control (MPC) can solve challenging tasks [2]. However, MPC often relies on constrained optimization, requiring constraint simplifications, precise modeling, and ad-hoc solutions to handle discontinuous dynamics in contact-rich tasks [3, 4]. While utilizing motion memory for warm-starting optimization can enhance performance [5, 6], the above limitations still persist.
Recently, Model Predictive Path Integral (MPPI) control [7] and its information-theoretic counterpart [8] addressed optimal control problems via importance sampling, mitigating challenges tied to constrained optimization algorithms dealing with non-convex constraints and discontinuous dynamics. However, substantial modeling remains necessary. In this paper, we propose a training-free model-based framework for real-time control of complex systems, where one designs only a task cost function, not the problem’s dynamics and contact models. We introduce the idea of using a general GPU-parallelizable physics simulator, IsaacGym [9], as the dynamic model for MPPI. By doing so, we create a robust framework that can generalize to various tasks. An overview is given in Fig. 1.
@@ I-A Related work
In contrast, Model Predictive Path Integral (MPPI) control [7, 8] is a sampling-based MPC approach that approximates optimal control via parallel sampling of input sequences. MPPI is gradient-free and well-suited for systems with non-linear, non-convex, discontinuous dynamics and cost functions. It has successfully controlled high-degree-of-freedom manipulators in real-time [16], incorporating self-collision avoidance using trained neural networks and collision-checking functions [17]. However, these approaches have limited interaction with the environment. In [18], the authors propose ensemble MPPI, a variation that handles complex tasks and adapts to parameter uncertainty, but the task modeling remains unclear, and no open-source implementation is available.
@@
A high number of samples is particularly crucial in tasks such as non-prehensile manipulation with robot manipulators. Traditional approaches often involve sampling end-effector trajectories on a plane, relying on additional controllers for robot actuation and learned models for predictions [22, 23]. For instance, Arruda et al. [22] use a forward-learned model trained on 326 real robot pushes. This model is employed by an MPPI controller to plan push manipulations as end-effector trajectories. Cong et al. [23] train a Long Short-Term Memory-based model to capture push dynamics using a dataset of 300 randomized objects. End-effector trajectories are sampled within a rectangular 2D workspace. Both methods require a separate controller to convert cartesian motions into joint commands, and both perform push manipulation through a sequence of pushes, resulting in discontinuous motion. These methods are not easily transferable to other robots, particularly for non-holonomic mobile pushing.
@@ I-B Contributions
In this paper, we present a novel combination of Model Predictive Path Integral (MPPI) control with the GPU-parallelizable simulator IsaacGym to tackle a variety of contact-rich motion planning problems. While our method utilizes existing techniques, it originally integrates them. The two key contributions of this work are:
  • The integration of the MPPI controller with IsaacGym, distinguishing our approach from prior works that used explicit dynamic models in MPPI. Our method facilitates collision checking and contact-rich manipulation tasks without requiring explicit modeling or learning of rigid body interactions. Our solution allows smooth real-time control of real-world high-degrees of freedom systems efficiently computing hundreds of rollouts in parallel.
  • A method that can address various motion planning challenges, including collision avoidance, prehensile and non-prehensile manipulation, and whole-body control with diverse robots. We provide an open-source implementation that can be reused and extended for various robotic tasks.
We demonstrate our method in several contact-rich problems with different robotic platforms and real-world experiments, as well as comparisons against specialized baselines.
@@ II-A Background theory on MPPI
Equation 3 and 4 demonstrate the approach taken to approximate the optimal control. Equation 4 represents a weighted average of sampled control inputs, while 3 assigns exponentially higher weights to less costly inputs. The first input $u_{0}^{*}$ of the sequence $U^{*}$ is applied to the system. Then the process is repeated.
@@ II-B Proposed algorithm
We now describe how we use MPPI with IsaacGym, summarized in Algorithm 1. We initialize an input sequence $U_{init}$ as a vector of zeroes with a length of $T$, where $T$ is the time horizon in steps. We then sample $K$ sequences of additive input noise $\mathcal{E}_{k}$ for exploring the input space around $U_{init}$. The key concept is that, instead of explicitly defining a nonlinear transition function ${f}$, we use IsaacGym to compute the next state ${x}_{t+1}$ given ${x}_{t}$ and control input ${v_{t}}$. This is done by reading the current state of the environment, resetting the state of the simulator to the observed values, and then applying the noisy control input sequence to simulate the state trajectories in IsaacGym. Note that these $K$ state trajectories can be computed independently of each other. We make use of this property to forward simulate all the rollouts in parallel leveraging the parallelization capabilities of IsaacGym. Instead of sampling from a Gaussian distribution, we follow the strategy of a recent paper [16] that proposes to sample Halton Splines instead for better exploration and smoother trajectories. Similar to [16], we calculate knot points as a Halton sequence. Given the knots, we fit a smooth spline approximation using standard Python modules and then we evaluate the spline at regular intervals to determine $\mathcal{E}_{k}$. Unlike [16], we do not update the variance of the sampling distribution. Instead, we keep it as a tuning parameter, constant during execution. Updating the variance as in [16] can lead to better convergence to a goal, but it also leads to stagnation of the control over time, which is harmful in the contact-rich tasks considered in this paper. Once the task begins, we reset our $K$ simulation environments on IsaacGym to the current observed world state $x$. In parallel, we can now roll out the sampled input sequences ${V}_{k}$ into state trajectories $Q_{k}$ using $K$ simulation environments on IsaacGym and compute their corresponding cost $S_{k}$ using the designed cost function $C$. The cost is discounted over the planning horizon $T$ by a factor $\gamma$ [16]:
@@
Next, we can compute the importance sampling weights $w_{k}$ as in 3. Note that the normalization factor $\eta$ is a useful metric to monitor, as it indicates the number of samples assigned significant weights. We use this insight to automatically tune the parameter $\beta$ to maintain $\eta$ within an upper and lower bound:
$\beta_{t+1}=\begin{cases}0.9\beta&\text{if}\ \eta>\eta_{max}\\ 1.2\beta&\text{if}\ \eta<\eta_{min}\\ {\color[rgb]{0,0,0}\beta}&{\color[rgb]{0,0,0}\text{otherwise}}\end{cases}$ (6)
@@
[…]
@@
[…]
@@ II-C2 Tackling model uncertainty
IsaacGym is designed to easily support domain randomization. We use this feature to randomize the object properties in each environment in case of contact-rich tasks, such that uncertainty is incorporated in every rollout for the MPPI. Effectively, this allows to account for uncertainty in environment perception. Specifically, starting from nominal physics properties, in every rollout objects are spawned with uncertainty on mass and friction nominal values, sampled from a uniform distribution. Additionally, the object size is also randomized with additive Gaussian noise, see Section III for experiment-specific details. Therefore, every simulation is different from the others, and all simulations are different from the world such that we can account for model mismatch. In a sense, we perform a sort of domain randomization in real-time to address the challenge of model uncertainty and imperfect perception. In the next section, we evaluate the performance of our method in different scenarios.
@@ III Experiments
We perform several experiments in three different categories: 1) motion planning and collision avoidance, 2) whole-body control of high DOF systems in contact-rich settings, and 3) non-prehensile manipulation. Experiments and simulations are carried out on an Alienware Laptop with Nvidia 3070 Ti graphics card. The software implementation consists of our open-source Python package that can easily be installed, tested, and extended to new robots and tasks. For the real-world tests, we used a Robot Operating System (ROS) wrapper to connect the robot to the planner, and we used a motion capture setup to determine the pose of manipulated objects.
@@ III-A Motion planning and collision avoidance
The compared methods show minimal differences in path length clearance for both examples (Fig. 3). However, our method consistently achieves faster convergence to the goal (Fig. 3 and 4, and Table I). This is attributed to the perfect representation of the robot’s collision shapes used in our method, compared to the enclosing spheres in the ForcesPro MPC and optimization fabrics. It should be noted that our approach incurs higher computational times (Table I) due to the physics simulations performed by IsaacGym. Despite this, our method remains competitive in motion planning applications and offers significant advantages in contact-rich tasks, as demonstrated in the following sections.
@@
Average
Point
1.0ms
2.5ms
55ms
Solver Time
Panda
1.4ms
51ms
63ms
Average
Point
7.4s
6.1s
2.7s
Time to Goal
Panda
9.6s
4.2s
0.8s
@@ III-B Prehensile manipulation with whole-body control
Figure 5: Whole-body motion of a mobile manipulator moving a cube from initial to desired location. $\omega_{t}=8$, $\omega_{O_{p}}=2$, $\omega_{ee}=1$, $\omega_{c}=0.1$, $T=6$, $dt=0.04$, $K=500$.
@@
$C_{pick}=C_{dist}+C_{ee}+C_{coll},$ (8)
where we only consider the Euclidean distance of the end-effector to the object, and the object to the goal:
$C_{dist}=\omega_{t}||p_{EE}-p_{O}||+\omega_{O_{p}}||p_{G}-p_{O}||,$
and we give the incentive to keep the end-effector pointing downwards, using pitch $\theta$ and roll $\phi$, with $C_{ee}=\omega_{ee}||[\phi,\theta]-[0,0]||$. By sampling all the DOF at once, including the base and the gripper, we achieve a fluid motion from start to end with no added heuristics for pick positions. We performed 5 pick-and-deliver tasks, and the time taken was 8.19 $\pm$ 1.51s. Mass and friction are randomized with 30% uncertainty sampled uniformly. Object size is randomized with additive Gaussian noise with a standard deviation of 1mm.
For smooth whole-body motions of high DOF systems, like this one, one requires a high number of samples (i.e. a few hundred). Empirically, when the number of samples is greater than 50, a GPU pipeline is computationally cheaper than a CPU and scales better. By using IsaacGym, we can compute all the 500 samples required for mobile manipulation in parallel, computing the next control input online at $25Hz$.
@@ III-C Non-prehensile manipulation
One advantage of using a physics simulator is that one can leverage generic physics rules for contacts, thus eliminating the need for learning or engineering specialized contact models. We demonstrate this in non-prehensile manipulation tasks involving a 7-DOF arm (Fig. 6) and two different mobile robots (Fig. 7, LABEL:, 8, LABEL: and 9). In Section III-C1, we apply our method to the two different tasks tackled in [22, 23] and we compare with their final results. Additionally, in Section III-C2, we demonstrate the ease of transferring our approach to different robots, including differential-drive.
@@ III-C1 Comparison with baselines for pushing with a robot arm
We consider two baselines for non-prehensile pushing. In the first one, [22] tackles the problem of pushing a relatively small object to a target pose with either 00 (Pose 1) or $90\deg$ (Pose 2). They also consider sequences of push actions starting far from the object. In the second baseline, [23] considers 5 relatively big objects and assumes the robot’s end effector is close to the object during execution. Since we do not have access to the same hardware, and the authors of the considered baselines do not provide their models and data, we only compare against their final results. Whenever their data cannot be retrieved for a certain result, we indicate this as n/a. We set up our simulation to match as close as possible the tasks in the baseline using the available information from the papers. Finally, we tune our method for the two tasks separately for a fair comparison with the individual baselines. The approach in [22] utilizes an MPPI in combination with a learned model for predicting pushing effects on an object. The authors sample 2D end-effector trajectories and then rely on inverse kinematic solvers, achieving push manipulation as a sequence of disconnected pushes. In contrast, we use MPPI to sample the control input directly as joint velocities in IsaacGym. By doing so, we achieve smooth continuous pushes where end-effector repositioning emerges naturally, and learning is not required. The cost function to be minimized for the task is:
@@
where $p_{G}$ and $\psi_{G}$ are the goal’s position and orientation, while $p_{R}$ and $p_{O}$ denote the end-effector tip and block positions, respectively. The cost function $C_{push\ align}$ promotes keeping the object between the robot and the goal. It is computed as $cos(\alpha)+1$, where $\alpha$ is the angle between the robot-object $(p_{R}-p_{O})$ and goal-object $(p_{G}-p_{O})$ vectors:
@@
We perform the same task as in [22] and compare the final results of pushing a squared object on a table surface to two poses (Pose 1 and 2) with a robot arm equipped with a stick. In Table II we report our findings, showing how our method can complete the task in a fraction of the time and with double the accuracy. We used the same evaluation metric of [22] for the final cost that is a weighted average of position and orientation errors: $1.5(|p_{R_{x}}-p_{O_{x}}|+|p_{R_{y}}-p_{O_{y}}|)+0.01|\psi_{O}-\psi_{G}|$. For every run, the manipulated object is also randomized in the same way as the rollouts. See the accompanying video to appreciate the actual behavior.
@@
Approach Start pose Final cost Time [s]
Baseline [22]
Pose 1
0.057
$\sim$ 240
Pose 2
0.079
n/a
Ours (sim)
Pose 1 0.029 ±$\pm 0.09$ 7.65 ±$\pm 2.62$
Pose 2 0.03 ±$\pm 0.12$ 8.38 ±$\pm 2.56$
We further compare our approach with [23] in terms of the success rate of non-prehensile manipulation. Particularly, we consider the same task settings, that is pushing 5 different objects to 3 different goal poses. To do so we simply change the objects in the simulation and slightly re-tune the MPPI [^2] Tuning: $\omega_{t}=5$, $\omega_{O_{p}}=25$, $\omega_{O_{r}}=21$, $\omega_{ee_{h}}=30$, $\omega_{ee_{r}}=0.3$, $\omega_{a}=45$, $dt=0.04$, $T=8$, $K=500$..
Since the trained models from [23] are not provided, we only compare the final results, reported in Table III. We performed 10 pushes per object, totaling 150 pushes. For the non-prehensile manipulation task with the robot arm, the mass and friction of manipulated objects have 30% uncertainty, and table friction has 90% uncertainty on the nominal value, sampled uniformly. Size is randomized with zero-mean additive Gaussian noise with a 2 mm standard deviation.
@@
Approach Metric
A
B
C
D
E
Baseline [23]
Success [%]
93.5
90.9
93.9
91.6
89.5
Time [s]
n/a
$\sim$ 24
$\sim$ 24
$\sim$ 24
$\sim$ 24
Ours (sim)
Success [%] 100 93.3 96.7 100
66.7
Time [s]
3.17 ±$\pm 1.81$
3.52 ±$\pm 1.84$
2.62 ±$\pm 1.03$
2.46 ±$\pm 0.47$
2.68 ±$\pm 1.51$
By using our method, we outperformed both baselines in terms of time to completion, accuracy, and success rate, except for one manipulated object. We achieve so without limiting the sampling to 2D end-effector trajectories, without the need for learned models, and without requiring inverse kinematics solvers.

### v3

@@ Abstract
We present a sampling-based model predictive control method that uses a generic physics simulator as the dynamical model. In particular, we propose a Model Predictive Path Integral controller (MPPI) that employs the GPU-parallelizable IsaacGym simulator to compute the forward dynamics of the robot and environment. Since the simulator implicitly defines the dynamic model, our method is readily extendable to different objects and robots, allowing one to solve complex navigation and contact-rich tasks. We demonstrate the effectiveness of this method in several simulated and real-world settings, including mobile navigation with collision avoidance, non-prehensile manipulation, and whole-body control for high-dimensional configuration spaces. This is a powerful and accessible open-source tool to solve many contact-rich motion planning tasks. Code and videos: https://autonomousrobots.nl/paper_websites/isaac-mppi
@@ I Introduction
As robots become increasingly integrated into our daily lives, their ability to navigate and interact with the environment is becoming more important than ever. From collision avoidance to moving obstacles out of the way to pick up some objects, robots must be able to plan their motions while accounting for contact with their surroundings. At the same time, robotic platforms require many Degrees Of Freedom (DOF) to achieve agile and dexterous movements. All this poses many challenging problems to motion planners, such as collision-free navigation in complex and dynamic environments, high DOF mobile manipulation, contact-rich tasks such as picking and pushing, and in-hand manipulation. Solutions to these challenges exist but are often specialized and not easily transferable to different scenarios. Learning-based approaches, for example, can leverage physics simulators to train policies for complex tasks but require extensive training and resources. For instance, [1] took years to develop, utilizing 6144 CPU cores and 50 hours of training to learn a policy for in-hand cube manipulation.
Figure 1: Scheme of the proposed method using IsaacGym as the dynamic model for MPPI. At each time step, IsaacGym is reset to the current world’s state $x$, and random input sequences ${V}$ are applied for the horizon $T$, to every environment. MPPI uses the resulting rolled-out trajectories to approximate the optimal control $u_{0}^{*}$ given a cost function $C$.
On the other hand, model-based approaches like Model Predictive Control (MPC) can solve challenging tasks [2]. However, MPC often relies on constrained optimization, requiring constraint simplifications, precise modeling, and ad-hoc solutions to handle discontinuous dynamics in contact-rich tasks [3, 4]. While utilizing motion memory for warm-starting optimization can enhance performance [5, 6], the above limitations still persist. Recently, Model Predictive Path Integral (MPPI) control [7] and its information-theoretic counterpart [8] addressed optimal control problems via importance sampling, mitigating challenges tied to constrained optimization algorithms dealing with non-convex constraints and discontinuous dynamics. However, substantial modeling remains necessary.
In this paper, we propose a training-free model-based framework for real-time control of complex systems, where one designs only a cost function, not the problem’s dynamics and contact models. We introduce the idea of using a general GPU-parallelizable physics simulator, IsaacGym [9], as the dynamic model for MPPI. This creates a robust framework that generalizes to various tasks. An overview is given in Fig. 1.
@@ I-A Related work
In contrast, Model Predictive Path Integral (MPPI) control [7, 8] is a sampling-based MPC approach that approximates optimal control via parallel sampling of input sequences. MPPI is gradient-free and well-suited for systems with non-linear, non-convex, discontinuous dynamics and cost functions. It has successfully controlled high-degree-of-freedom manipulators in real-time [16], incorporating self-collision avoidance using trained neural networks and collision-checking functions [17]. However, these approaches have limited interaction with the environment. In [18], the authors propose ensemble MPPI, a variation that handles complex tasks and adapts to parameter uncertainty. Still, the task modeling remains unclear, and no open-source implementation is available.
@@
A high number of samples is particularly crucial in tasks such as non-prehensile manipulation with robot manipulators. Traditional approaches often involve sampling end-effector trajectories on a plane, relying on additional controllers for robot actuation and learned models for predictions [22, 23]. For instance, Arruda et al. [22] use a forward-learned model trained on 326 real robot pushes. This model is employed by an MPPI controller to plan push manipulations as end-effector trajectories. Cong et al. [23] train a Long Short-Term Memory-based model to capture push dynamics using a dataset of 300 randomized objects. End-effector trajectories are sampled within a rectangular 2D workspace. Both methods require a separate controller to convert cartesian motions into joint commands, and both perform push manipulation through a sequence of pushes, resulting in discontinuous motion. These methods are not easily transferable to other robots, particularly non-holonomic mobile pushing.
@@ I-B Contributions
This paper presents a novel open-source implementation of Model Predictive Path Integral (MPPI) control with a generic physics simulator as the dynamical model. This enables the method to solve many contact-rich motion planning problems. The two key contributions of this work are:
  • The integration of the MPPI controller with the GPU-parallelizable simulator IsaacGym, distinguishing our approach from prior works in MPPI. Our method facilitates collision checking and contact-rich manipulation tasks leveraging the contact models and rigid body interactions included in the simulator without requiring gradients. Our solution allows smooth real-time control of real-world systems with high degrees of freedom, efficiently computing hundreds of rollouts in parallel.
  • A versatile method applicable to various motion planning challenges, including collision avoidance, prehensile and non-prehensile manipulation, and whole-body control with diverse robots. We provide an open-source implementation that can be readily reused and extended to heterogeneous robots and tasks.
We perform many contact-rich tasks with several robotic platforms and real-world experiments. We include omnidirectional and differential drive robots and fixed or mobile manipulators and compare against many specialized baselines.
@@ II-A Background theory on MPPI
3 and 4 demonstrate the approach taken to approximate the optimal control. 4 represents a weighted average of sampled control inputs, while 3 assigns exponentially higher weights to less costly inputs. The first input $u_{0}^{*}$ of the sequence $U^{*}$ is applied to the system. Then the process is repeated.
@@ II-B Proposed algorithm
We now describe how we use MPPI with IsaacGym, summarized in Algorithm 1. We initialize an input sequence $U_{init}$ as a vector of zeroes with a length of $T$, where $T$ is the time horizon in steps. We then sample $K$ sequences of additive input noise $\mathcal{E}_{k}$ for exploring the input space around $U_{init}$. The key concept is that, instead of explicitly defining a nonlinear transition function ${f}$, we use IsaacGym to compute the next state ${x}_{t+1}$ given ${x}_{t}$ and control input ${v_{t}}$. This is done by reading the current state of the environment, resetting the state of the simulator to the observed values, and then applying the noisy control input sequence to simulate the state trajectories in IsaacGym. Note that these $K$ state trajectories can be computed independently of each other. We use this property to forward and simulate all the rollouts in parallel, leveraging the parallelization capabilities of IsaacGym. Instead of sampling from a Gaussian distribution, we follow the strategy of a recent paper [16] that proposes to sample Halton Splines instead for better exploration and smoother trajectories. Similar to [16], we fit B-Splines to inputs sampled from a Halton sequence using standard Python modules and then we evaluate the spline at regular intervals to retrieve $\mathcal{E}_{k}$. Unlike [16], we do not update the variance of the sampling distribution. Instead, we keep it as a tuning parameter, constant during execution. Updating the variance as in [16] can lead to better convergence to a goal, but it also leads to stagnation of the control over time, which is harmful in the contact-rich tasks considered in this paper. Once the task begins, we reset our $K$ simulation environments on IsaacGym to the current observed world state $x$. In parallel, we can now roll out the sampled input sequences ${V}_{k}$ into state trajectories $Q_{k}$ using $K$ simulation environments on IsaacGym and compute their corresponding cost $S_{k}$ using the designed cost function $C$. The cost is discounted over the planning horizon $T$ by a factor $\gamma$ [16]:
@@
Next, we can compute the importance sampling weights $w_{k}$ as in 3. The normalization factor $\eta$ is a useful metric to monitor, as it indicates the number of samples assigned significant weights. We use this to tune $\beta$ for the next iteration such that $\eta$ is maintained within an upper and lower bound:
$\beta_{t+1}=\begin{cases}0.9\beta_{t}&\text{if}\ \eta>\eta_{max}\\ 1.2\beta_{t}&\text{if}\ \eta<\eta_{min}\\ \beta_{t}&\text{otherwise}\end{cases}$ (6)
@@
10:     $\beta\leftarrow\textit{updateBeta}(\beta,\eta)$triangleright 6
@@
12:     $U_{init}\leftarrow\textit{timeShift}(U^{*})$
13:     $\textit{applyInput}(u_{0}^{*})$
@@ II-C2 Tackling model uncertainty
IsaacGym is designed to easily support domain randomization. We use this feature to randomize the object properties in each environment in case of contact-rich tasks, such that uncertainty is incorporated in every rollout for the MPPI. Effectively, this allows to account for uncertainty in environment perception. Specifically, starting from nominal physics properties, in every rollout objects are spawned with uncertainty on mass and friction nominal values, sampled from a uniform distribution. Additionally, the object size is also randomized with additive Gaussian noise, see Section III for experiment-specific details. Therefore, every simulation is different from the others, and all simulations are different from the world such that we can account for model mismatch. In a sense, we perform a sort of domain randomization in real-time to address the challenge of model uncertainty and imperfect perception.
@@ III Experiments
We perform several experiments in three different categories: 1) motion planning and collision avoidance, 2) whole-body control of high DOF systems in contact-rich settings, and 3) non-prehensile manipulation. Experiments and simulations are conducted on an Alienware Laptop with Nvidia 3070 Ti graphics card. The software implementation consists of our open-source Python package that can easily be installed, tested, and extended to new robots and tasks. In real-world tests, we used a Robot Operating System (ROS) wrapper to connect the robot to the planner and a motion capture system to determine the pose of manipulated objects. Our implementation allows for position, velocity, and torque control. In this paper, all robots are velocity-controlled except for the mobile manipulator in Section III-B, which is torque-controlled.
@@ III-A Motion planning and collision avoidance
The compared methods show minimal differences in path length clearance for both examples (Fig. 3). However, our method consistently reaches the goal faster (Fig. 3 and 4, and Table I). This is attributed to the perfect representation of the robot’s collision shapes used in our method, compared to the enclosing spheres in the ForcesPro MPC and optimization fabrics. It should be noted that our approach incurs higher computational times (Table I) due to the physics simulations performed by IsaacGym. Despite this, our method remains competitive in motion planning applications and offers significant advantages in contact-rich tasks, as demonstrated in the following sections.
@@
Average Point 1.0ms 2.5ms 55ms
Solver Time Panda 1.4ms 51ms 63ms
Average Point 7.4s 6.1s 2.7s
Time to Goal Panda 9.6s 4.2s 0.8s
@@ III-B Prehensile manipulation with whole-body control
Figure 5: Whole-body motion of a mobile manipulator moving a cube from initial to a desired location.
@@
$C_{pick}=C_{dist}+C_{pose}+C_{coll}+C_{vel},$ (8)
where we consider the Euclidean distance of the end-effector to the object and the object to the goal: $C_{dist}=\omega_{t}||p_{EE}-p_{O}||+\omega_{O_{p}}||p_{G}-p_{O}||$. We give an incentive to keep the robot in a comfortable pose by penalizing deviations from a desired arm and gripper pose, end-effector orientation, as well as imposing a minimum end-effector height $C_{pose}=C_{Parm}+C_{Pgrip}+C_{Oee}+C_{Hee}$. We minimize collisions penalizing the forces on the table $C_{coll}$. Lastly, we penalize high arm and base velocities $C_{vel}=C_{Varm}+C_{Vbase}$ since, in this experiment, we torque-control the robot. By sampling all the DOF at once, including the base and the gripper, we achieve a fluid motion from start to end with no added heuristics for pick positions. We performed ten pick-and-deliver tasks, and the time taken was 15.67 $\pm$ 7.21s. The high standard deviation is because sometimes the cube falls, but the robot can recover by picking it up again from the floor.
For smooth whole-body motions of high DOF systems like this, many samples are required. Empirically, when the number of samples exceeds 50, a GPU pipeline is computationally cheaper than a CPU and scales better. Using IsaacGym, we can compute all the 750 samples required for mobile manipulation in parallel, computing the next control input online at $25Hz$.
@@ III-C Non-prehensile manipulation
One advantage of using a physics simulator is that one can leverage generic physics rules for contacts, thus eliminating the need for learning or engineering specialized contact models. We demonstrate this in non-prehensile manipulation tasks involving a 7-DOF arm (Fig. 6) and two different mobile robots (Fig. 7, LABEL:, 8, LABEL: and 9). In Section III-C1, we apply our method to the two pushing tasks tackled in [22, 23], and we compare with their final results. Additionally, in Section III-C2, we demonstrate the ease of transferring our approach to different robots, including differential-drive.
@@ III-C1 Comparison with baselines for pushing with a robot arm
We consider two baselines for non-prehensile pushing. In the first one, [22] tackles the problem of pushing a relatively small object to a target pose with either 00 (Pose 1) or $90\deg$ (Pose 2). They also consider sequences of push actions starting far from the object. In the second baseline, [23] considers 5 relatively big objects and assumes the robot’s end effector is close to the object during execution. Since we do not have access to the same hardware, and the authors of the considered baselines do not provide their models and data, we only compare against their final results. We set up our simulation to match as close as possible the tasks in the baseline using the available information from the papers. Finally, we tune our method for the two tasks separately for a fair comparison with the individual baselines. The approach in [22] utilizes an MPPI in combination with a learned model for predicting pushing effects on an object. The authors sample 2D end-effector trajectories and then rely on inverse kinematic solvers, achieving push manipulation as a sequence of disconnected pushes. In contrast, we use MPPI to sample the control input directly as joint velocities in IsaacGym. By doing so, we achieve smooth continuous pushes where end-effector repositioning emerges naturally, and learning is not required. The cost function to be minimized for the task is:
@@
where $p_{G}$ and $\psi_{G}$ are the goal’s position and orientation, while $p_{R}$ and $p_{O}$ denote the end-effector tip and block positions, respectively. The cost function $C_{push\ align}$ promotes keeping the object between the robot and the goal. It is computed as $cos(\alpha)+1$, where $\alpha$ is the angle between the robot-object $(p_{R}-p_{O})$ and goal-object $(p_{G}-p_{O})$ vectors and $+1$ is added to make the cost term always non-negative [26]:
@@
We perform the same task as in [22] and compare the final results of pushing a squared object on a table surface to two poses (Pose 1 and 2) with a robot arm equipped with a stick. In Table II, we report our findings, with our method showing double the accuracy. Our approach performs continuous pushes, unlike the baseline that stops for replanning after each short push. Thus, we complete either task in approximately 8 seconds, while the baseline takes approximately 4 minutes. We used the same evaluation metric of [22] for the final cost that is a weighted average of position and orientation errors: $1.5(|p_{G_{x}}-p_{O_{x}}|+|p_{G_{y}}-p_{O_{y}}|)+0.01|\psi_{O}-\psi_{G}|$. For every run, the object is also randomized in the same way as the rollouts. See the accompanying video for the actual behavior.
@@
Approach Start pose Final cost
Baseline [22] Pose 1 0.057
Pose 2 0.079
Ours (sim) Pose 1 0.029 ±$\pm 0.09$
Pose 2 0.03 ±$\pm 0.12$
We further compare our approach with [23] in terms of the success rate of non-prehensile manipulation. Particularly, we consider the same task settings, pushing 5 different objects to 3 different goal poses. To do so we simply change the objects in the simulation and slightly re-tune the MPPI [^2] Tuning: $\omega_{t}=5$, $\omega_{O_{p}}=25$, $\omega_{O_{r}}=21$, $\omega_{ee_{h}}=30$, $\omega_{ee_{r}}=0.3$, $\omega_{a}=45$, $dt=0.04$, $T=8$, $K=500$..
Since the trained models from [23] are not provided, we only compare the final results, reported in Table III. Again, thanks to the continuous pushes, our method takes about 3 seconds per task, while the baseline needs about 24 seconds. We performed 10 pushes per object, totaling 150 pushes. For the non-prehensile manipulation task with the robot arm, the mass and friction of manipulated objects have 30% uncertainty, and table friction has 90% uncertainty on the nominal value, sampled uniformly. Size is randomized with zero-mean additive Gaussian noise with a 2 mm standard deviation.
@@
Approach Metric A B C D E
Baseline [23] Success [%] 93.5 90.9 93.9 91.6 89.5
Ours (sim) Success [%] 100 93.3 96.7 100 66.7
Our method outperforms both baselines in terms of time to completion, accuracy, and success rate, except for one manipulated object. We achieve this without limiting the sampling to 2D end-effector trajectories, without needing learned models, and without requiring inverse kinematics solvers.